A 5.0-kg mass is suspended by a string from the ceiling of an elevator that is moving upward with a speed which is decreasing at a constant rate of 2.0 m/s in eachsecond. What is the tension in the string supporting the mass?
Answer #1

bybalancing forces

mg - T = ma

T = 5g - 5(2) = 39.05 N

answered by: Ronasia Mcbride
Answer #2
T=m(g-a)=5(9.8-2)=39N
answered by: Bliss